# From Alignment to Advancement: Bootstrapping Audio-Language Alignment with Synthetic Data

### v2

@@ VII Conclusion
In this work, we introduce BALSa, a framework that synthesizes general-purpose caption-style alignment data for training ALLM. This framework reduces the reliance on large collections of task-specific QA datasets. Our experimental results demonstrate that the proposed approach achieves comparable or superior performance to previous approaches while improving data and computational efficiency. BALSa effectively mitigates audio hallucinations, further reinforcing its robustness in real-world applications. Beyond single-audio alignment, we extend to more complex audio understanding settings. In multi-audio scenarios, this extension further enhances modality alignment and leads to better performance on audio understanding and reasoning tasks. We further examine the role of discrimination-based training and observe that learning to handle multiple audio inputs alone is sufficient to improve overall understanding, without requiring explicit comparative objectives. Overall, this paper provides a scalable and efficient pathway for developing and advancing audio-aware large language models.

### v3

@@ VII Conclusion
In this work, we introduce BALSa, a efficient framework that synthesizes general-purpose alignment data for training Audio-Aware Large Language Models (ALLMs). A key innovation of BALSa is applying a synthetic data generation paradigm, previously focused on speech, to the general audio domain, thereby reducing the reliance on large collections of task-specific QA datasets. Our experimental results demonstrate that the proposed approach achieves competitive performance against state-of-the-art models with remarkable data efficiency, using only a fraction of the training data of strong baselines. Crucially, by incorporating a contrastive-style strategy with synthetic negative samples, BALSa effectively mitigates audio hallucinations, enhancing the model’s reliability in real-world applications. Beyond single-audio alignment, we extend the framework to multi-audio scenarios with BALSa-MA. This extension further enhances modality alignment and leads to substantial performance gains on complex audio understanding and reasoning tasks. Our analyses reveal that learning to process multiple audio inputs simultaneously is the key driver for this improvement, even without explicit comparative objectives. Finally, we verified the framework’s robustness and scalability across different backbone models and fine-tuning strategies. Overall, this paper provides a practical, scalable, and efficient pathway for developing robust and reliable audio-aware large language models.